A sample of 5 months of sales data provided the following information:


1.

A sample of 5 months of sales data provided the following information: 

Month 1 2 3 4 5

Units Sold  95 120 82 95 93

a. Develop a point estimate of the population mean number of units sold per month (to 1 decimal)  

b. Develop a point estimate of the population standard deviation (to 2 decimals).


2.

The following data are from a simple random sample. 

5, 8, 13, 6, 13, 15 

a. What is the point estimate of the population mean? 

b. What is the point estimate of the population standard deviation (to 1 decimal)? 
Answer #1

(first part) population mean=sum(x)/n=485/5=97

population standard deviation=sqrt(sum((x-mean)2)/n)=sqrt(778/5)=12.47

Month units sold (x) x-mean (x-mean)2
1 95 -2 4
2 120 23 529
3 82 -15 225
4 95 -2 4
5 93 -4 16
n= 5 6 7
sum= 485 0 778
mean=sum/n= 485/5=97

(second part)

population mean=sum(x)/n=60/6=10

population standard deviation=sqrt(sum((x-mean)2)/n)=sqrt(88/6)=3.83

Month units sold (x) x-mean (x-mean)2
1 5 -5 25
2 8 -2 4
3 13 3 9
4 6 -4 16
5 13 3 9
6 15 5 25
n= 6 6 6
sum= 60 0 88
mean=sum/n= 60/6=10